# Exploiting Multimodal Synthetic Data for Egocentric Human-Object Interaction Detection in an Industrial Scenario

### v1

@@ 6.5.1 Pre-training on 100 Days Of Hands
Table 6: EHOI detection results on EgoISM-HOI-Real test data of models pre-trained on 100DOH dataset.
@@
Using synthetic and real-world training data (row 3) leads to the best or second-best results for all the evaluation metrics. In particular, the improvements over the model which uses only EgoISM-HOI-Real data (rows 2) are significant in the metrics affected by the active objects with a +20,44% (38.54 vs 18.10) for the mAP Hand+Obj and +19,68 (37.37 vs 17.69) for the mAP Hand+All. Considering the mAP Hand+All metric, it is worth noting that the model trained only on EgoISM-HOI-Synth (row 1) surpasses the model trained on the EgoISM-HOI-Real training data (row 2) with an improvement of +5,5% (23.19 vs 17.69).
@@ 7 Conclusion
We studied egocentric human-object interactions in an industrial domain. Due to the expensiveness of collecting and labeling real in-domain data in the considered context, we proposed a pipeline and a tool that leverages 3D models of the objects and the considered environment to generate synthetic images of EHOIs automatically labeled and additional data signals, such as depth maps and instance segmentation masks. Exploiting our pipeline, we presented EgoISM-HOI, a new multimodal dataset of synthetic and real EHOI images in an industrial scenario with rich annotations of hands and objects. We investigated the potential of using multimodal synthetic data to pre-train an EHOI detection system and demonstrated that our proposed method outperforms class-agnostic baselines based on the state-of-the-art method of Shan et al. (2020). Future work will investigate how the knowledge inferred by our method can be valuable for other related tasks such as next active object detection or action recognition. To encourage research on the topic, we publicly released the datasets and the source code of the proposed system, together with pre-trained models, on our project web page: https://iplab.dmi.unict.it/egoism-hoi.

### v2

@@ 6.4.2 Comparison with VISOR baseline
Table 7: EHOI detection results on EgoISM-HOI-Real test data of models pre-trained on 100DOH and VISOR datasets. The grey row shows the results obtained by the model trained only on the EgoISM-HOI dataset, serving as the baseline for this set of experiments.
@@
VISOR EgoISM-HOI-Synth 90.58 89.24 36.07 24.78 24.49
VISOR EgoISM-HOI-Real 90.65 90.28 57.84 29.51 29.40
VISOR EgoISM-HOI 90.74 90.35 56.71 39.11 38.95
Table 8: Impact of pretraining with varying numbers of EgoISM-HOI-Synth images.
Pretraining Data #Pretraining Images Finetuning Data mAP H.+All
EgoISM-HOI-Synth 1,010 EgoISM-HOI-Real 32.83
EgoISM-HOI-Synth 2,020 EgoISM-HOI-Real 33.01
EgoISM-HOI-Synth 4,040 EgoISM-HOI-Real 34.96
EgoISM-HOI-Synth 10,100 EgoISM-HOI-Real 35.34
EgoISM-HOI-Synth 20,788 EgoISM-HOI-Real 35.47
6.5 Additional results
In this section, we show an additional set of experiments with the aim of 1) demonstrating how using domain-specific synthetic data improves the performance of a system pre-trained on out-of-domain large-scale datasets (Section 6.5.1), 2) analyzing the impact of varying quantities of synthetic images on the pre-training of the system (Section 6.5.2), 3) investigating the influence of the different modules within our system (Section 6.5.3), 4) showing the potential of using synthetic data for the related task of Object Detection (Section 6.5.4).
6.5.1 Pre-training on 100 Days Of Hands and VISOR
To further confirm the usefulness of synthetic data, we performed additional experiments where we pre-trained models on 100DOH (Shan et al., 2020) and VISOR (Darkhalil et al., 2022) datasets which were then fine-tuned considering the proposed EgoISM-HOI dataset. These experiments aim to demonstrate how leveraging domain-specific synthetic data enhances the performance of a system pre-trained on a large amount of out-of-domain real-world data. The results shown in Table 7 highlight that employing synthetic data in the fine-tuning phase consistently led to superior performance for both 100DOH and VISOR pre-trained models. Considering the 100DOH pre-trained model, the combination of synthetic data with real-world data (row 4) significantly enhances metrics influenced by active objects (i.e., mAP Hand+Obj and mAP Hand+All). Specifically, we observed an improvement of +20.44% (38.54% vs. 18.10%) for the mAP Hand+Obj and of +19.68% (37.37% vs. 17.69%) for the mAP Hand+All, compared to the model trained only on real-world data (row 3). Similar considerations can be made for the experiments performed considering the VISOR dataset as pre-training. Indeed, the model trained using both synthetic data with real-world data (row 7) outperforms their counterparts trained only on real-world data (row 6). In this last case, we observe an improvement of +9.60% (39.11% vs. 29.51%) for mAP Hand+Obj and of +9.05% (38.95% vs. 29.40%) for mAP Hand+All. It is important to note that, as seen in previous investigations (see Section 6.2), for metrics influenced exclusively by hands (for example, AP Hand, AP H+Side and AP H+State), we observed minor or no improvements when synthetic data were used, compared to models trained exclusively on real-world data. Lastly, it’s worth noting that the results obtained from the baseline model, i.e. the model trained only on the EgoISM-HOI data (first row), obtained comparable or even superior performance with respect to models trained using real-world data (i.e., 100DOH and VISOR). This further highlights the usefulness of synthetic data in enhancing HOI model performances.
6.5.2 Effect of Varying Synthetic Data Quantities on Pretraining
We conducted a series of experiments using a varying number of synthetic images from EgoISM-HOI-Synth during pre-training. Table 8 collects the results of this experiment. The results in terms of mAP Hand+All show an evident trend of improvement in performances as the number of pre-training images increases. Specifically, starting with 1,010 synthetic images, the model achieved an initial performance of 32.83%. Subsequently, by doubling the pre-training images to 2,020, a slight improvement was observed, reaching 33.01%. Performance further increased with 4,040 synthetic images (34.96%) and with 10,100 synthetic images (35.34%). Finally, using all the 20,788 synthetic images resulted in the highest performance with a mAP Hand+All of 35.47%. This highlights the impact of a larger synthetic dataset on enhancing the model’s performance.
6.5.3 Understanding the Weighting of Various Modalities in the System
Our proposed approach combines the predictions of the hand contact state produced by the hand side classifier and multimodal hand state classifier modules to produce the final hand contact state prediction. We have analyzed the impact of these modules in our framework by assigning a weight to the prediction of each module. Table 9 collect the results of these experiments in terms of AP Hand+State and mAP Hand+All evaluation measures. The obtained results show that as the weight shifts towards a balance between HS and MHS modules, the performances improve. This highlights the importance of both modules. Specifically, the system achieves its peak performance at 50% for both modules, obtaining 58.40% and 36.51% for the textitAP Hand+State and mAP Hand+All, respectively.
Table 9: The table demonstrates how varying the weighting between the hand side classifier (HS) and the multimodal hand state classifier (MHS) affects the performance of the system in terms of AP Hand+State and mAP Hand+All.
HS Weight MHS Weight AP H.+State mAP H.+All
10% 90% 57.05 35.39
20% 80% 57.35 35.66
30% 70% 57.66 36.08
40% 60% 57.96 36.22
50% 50% 58.40 36.51
60% 40% 58.35 35.98
70% 30% 58.09 36.38
80% 20% 57.89 35.86
90% 10% 57.70 35.84
@@ 7 Conclusion
We studied egocentric human-object interactions in an industrial domain. Due to the expensiveness of collecting and labeling real in-domain data in the considered context, we proposed a pipeline and a tool that leverages 3D models of the objects and the considered environment to generate synthetic images of EHOIs automatically labeled and additional data signals, such as depth maps and instance segmentation masks. Exploiting our pipeline, we presented EgoISM-HOI, a new multimodal dataset of synthetic and real EHOI images in an industrial scenario with rich annotations of hands and objects. We investigated the potential of using multimodal synthetic data to pre-train an EHOI detection system and demonstrated that our proposed method outperforms class-agnostic baselines based on the state-of-the-art method of Shan et al. (2020). Future work will investigate how the knowledge inferred by our method can be valuable for other related tasks such as next active object detection or action recognition. Additionally, there is a need to consider the problem of handling and accurately recognizing simultaneous interactions with multiple objects. To encourage research on the topic, we publicly released the datasets and the source code of the proposed system, together with pre-trained models, on our project web page: https://iplab.dmi.unict.it/egoism-hoi.